As a promotion, a certain car dealer gave away a free watch to anyone who visited his dealership and test drove a car. But 5% of the watches were defective.
a. Find the probability that among 50 such watches selected at random, exactly one is defective.
b. Find the probability that among 50 such watches selected at random, none is defective.
c. Find the probability that among 50 such watches selected at random, at least one is defective.

Answer by stanbon(75887)

As a promotion, a certain car dealer gave away a free watch to anyone who visited his dealership and test drove a car. But 5% of the watches were defective.
---
Binomial problems with n = 50 and p(defective) = 0.05
-------------------
a. Find the probability that among 50 such watches selected at random, exactly one is defective.
P(x=1) = 50C1*0.05*0.95^49 = 0.2025
============================================
b. Find the probability that among 50 such watches selected at random, none is defective.
P(x = 0) = 0.95^50 = 0.0759
============================================
c. Find the probability that among 50 such watches selected at random, at least one is defective.
P(x>=1) = 1-P(x=0) = 1-0.0759 = 0.9231
=========================================
Cheers,
Stan H.
============